# A deep backward regression-based scheme for high-dimensional nonlinear partial differential equations

### 16 Mar 2026 v1

@@ 3.2 Error analysis
$\mathbb{E}\left[\left|\mathcal{Y}_{i}-S_{y,i}\left(p_{i}^{m},X_{i+1}^{\pi,m}\right)\right|^{2}\right]=\mathbb{E}\left[\left|R_{i,2}^{y}\right|^{2}\right].$ (3.27)
@@
$\mathbb{E}\left[\left|S_{y,i}\left(p_{i}^{m},X_{i+1}^{\pi,m}\right)-\mathcal{U}^{m}_{t_{i}}\right|^{2}\right]$
@@
$\leq(1+Ch)\mathbb{E}\left[\left|\mathcal{NN}_{y,i+1,1+d,n}^{\varrho,L,\mathbf{n}}(p_{i+1}^{m},\theta_{y,i+1}^{*})-\mathbb{E}_{i}\left[\mathcal{NN}_{y,i+1,1+d,n}^{\varrho,L,\mathbf{n}}(p^{m}_{i+1},\theta^{*}_{y,i+1})\right]\right|^{2}\right]$
$\quad+Ch\mathbb{E}\left[\left|\mathcal{NN}_{z,i,1+d,n\times d}^{\varrho,L,\mathbf{n}}(p_{i}^{m},\theta_{z,i}^{*})-\mathbb{E}_{i}\left[\mathcal{V}^{m}_{t_{i}}\right]\right|^{2}\right]$
@@
$\leq(1+Ch)\mathbb{E}\left[\left|\mathcal{NN}_{y,i+1,1+d,n}^{\varrho,L,\mathbf{n}}(p^{m}_{i+1},\theta_{y,i+1}^{*})-\mathbb{E}_{i}\left[\mathcal{NN}_{y,i+1,1+d,n}^{\varrho,L,\mathbf{n}}(p^{m}_{i+1},\theta^{*}_{y,i+1})\right]\right|^{2}\right]$
$\quad+Ch\mathbb{E}\left[\left|Z_{i}^{\pi,m}-R_{i,2}^{z}-R_{i,1}^{z}-\mathcal{V}^{m}_{t_{i}}\right|^{2}\right]$
@@
$\leq(1+Ch)\mathbb{E}\left[\left|\mathcal{NN}_{y,i+1,1+d,n}^{\varrho,L,\mathbf{n}}(p^{m}_{i+1},\theta_{y,i+1}^{*})-\mathbb{E}_{i}\left[\mathcal{NN}_{y,i+1,1+d,n}^{\varrho,L,\mathbf{n}}(p^{m}_{i+1},\theta^{*}_{y,i+1})\right]\right|^{2}\right]$
$\quad+Ch\mathbb{E}\left[\left|R_{i,1}^{z}\right|^{2}\right]+Ch\mathbb{E}\left[\left|R_{i,2}^{z}\right|^{2}\right]$
$\quad+Ch\mathbb{E}\left[\left|\mathbb{E}_{i}\left[Y_{i+1}^{\pi,m}\frac{(\Delta W_{i}^{m})^{\top}}{h}\right]-\mathbb{E}_{i}\left[\mathcal{NN}_{y,i+1,1+d,n}^{\varrho,L,\mathbf{n}}(p^{m}_{i+1},\theta^{*}_{y,i+1})\frac{(\Delta W_{i}^{m})^{\top}}{h}\right]\right|^{2}\right]$
$\leq(1+Ch)\mathbb{E}\left[\left|\mathcal{NN}_{y,i+1,1+d,n}^{\varrho,L,\mathbf{n}}(p^{m}_{i+1},\theta_{y,i+1}^{*})-\mathbb{E}_{i}\left[\mathcal{NN}_{y,i+1,1+d,n}^{\varrho,L,\mathbf{n}}(p^{m}_{i+1},\theta^{*}_{y,i+1})\right]\right|^{2}\right]$
$\quad+Ch\mathbb{E}\left[\left|R_{i,1}^{z}\right|^{2}\right]+Ch\mathbb{E}\left[\left|R_{i,2}^{z}\right|^{2}\right]+C\mathbb{E}\left[\left|Y_{i+1}^{\pi,m}-\mathcal{NN}_{y,i+1,1+d,n}^{\varrho,L,\mathbf{n}}(p^{m}_{i+1},\theta^{*}_{y,i+1})\right|^{2}\right]$
$\leq C(1+h)\mathbb{E}\left[\left|\mathcal{NN}_{y,i+1,1+d,n}^{\varrho,L,\mathbf{n}}(p^{m}_{i+1},\theta_{y,i+1}^{*})\right|^{2}\right]$
$\quad+C\mathbb{E}\left[\left|R_{i+1,1}^{y}\right|^{2}+h\left|R_{i,1}^{z}\right|^{2}\right]+C\mathbb{E}\left[\left|R_{i+1,2}^{y}\right|^{2}+h\left|R_{i,2}^{z}\right|^{2}\right].$
From (3.8), we deduce
$\mathbb{E}\left[\left|S_{y,i}\left(p_{i}^{m},X_{i+1}^{\pi,m}\right)-\mathcal{U}^{m}_{t_{i}}\right|^{2}\right]$ (3.28)
$\leq C(1+h)\mathbb{E}\left[\left|Y_{i+1}^{\pi,m}-R_{i+1,2}^{y}-R_{i+1,1}^{y}\right|^{2}\right]+C\mathbb{E}\left[\left|R_{i+1,1}^{y}\right|^{2}+h\left|R_{i,1}^{z}\right|^{2}\right]+C\mathbb{E}\left[\left|R_{i+1,2}^{2}\right|^{2}+h\left|R_{i,2}^{z}\right|^{2}\right]$
$\leq C(1+h)\mathbb{E}\left[\left|Y_{i+1}^{\pi,m}-Y_{t_{i+1}}\right|^{2}+\left|Y_{t_{i+1}}\right|^{2}\right]+C\mathbb{E}\left[\left|R_{i+1,1}^{y}\right|^{2}+h\left|R_{i,1}^{z}\right|^{2}\right]+C\mathbb{E}\left[\left|R_{i+1,2}^{2}\right|^{2}+h\left|R_{i,2}^{z}\right|^{2}\right]$
$\leq C(1+h)+C\mathbb{E}\left[\left|R_{i+1,1}^{y}\right|^{2}+h\left|R_{i,1}^{z}\right|^{2}\right]+C\mathbb{E}\left[\left|R_{i+1,2}^{2}\right|^{2}+h\left|R_{i,2}^{z}\right|^{2}\right],$
where we use the estimates $\mathbb{E}\left[\left|Y_{i}^{\pi,m}-Y_{t_{i}}\right|^{2}\right]\leq Ch$ and $\mathbb{E}\left[\max\limits_{0\leq i\leq N}\left|Y_{t_{i}}\right|^{2}\right]<+\infty$ (see [19, 36, 28]) in the last inequalities. Plugging (3.27), (3.28) into (3.26), we derive
$\mathbb{E}\left[\left|\mathcal{Y}_{i}-\mathcal{U}_{t_{i}}\right|^{2}\right]\leq$ $C\mathbb{E}\left[\left|R_{i,2}^{y}\right|^{2}+\left|R_{i+1,2}^{y}\right|^{2}+h\left|R_{i,2}^{z}\right|^{2}+\left|R_{i+1,1}^{y}\right|^{2}+h\left|R_{i,1}^{z}\right|^{2}\right].$ (3.29)
Inserting (3.29) into (3.25), we deduce
$\max\limits_{0\leq i\leq N-1}\mathbb{E}\left[\left|Y_{t_{i}}-\mathcal{Y}_{i}\right|^{2}\right]$ $\leq C\mathbb{E}\left[\left|g(X_{T})-g(X_{N}^{\pi,m})\right|^{2}\right]+Ch+C\mathbb{E}\left[\sum\limits_{i=0}^{N-1}\int_{t_{i}}^{t_{i+1}}|Z_{s}-\bar{Z}_{t_{i}}|^{2}ds\right]$ (3.30)
$\quad+C\sum\limits_{i=0}^{N-1}\mathbb{E}\left[N\left|R_{i+1,2}^{y}\right|^{2}+\left|R_{i,2}^{z}\right|^{2}+N\left|R_{i+1,1}^{y}\right|^{2}+\left|R_{i,1}^{z}\right|^{2}\right].$
Step 3. Now, we prove the consistency of the $Z$-component. By (3.17) and (3.21), we obtain
$\mathbb{E}\left[\int_{t_{i}}^{t_{i+1}}|Z_{s}-\mathcal{V}^{m}_{t_{i}}|^{2}ds\right]$ (3.31)
@@
$\quad+4d\mathbb{E}\left[\left|Y_{t_{i+1}}-\mathcal{Y}_{i+1}\right|^{2}\right]+4d\mathbb{E}\left[\left|\mathcal{Y}_{i+1}-\mathcal{NN}_{y,i+1,1+d,n}^{\varrho,L,\mathbf{n}}(p^{m}_{i+1},\theta^{*}_{y,i+1})\right|^{2}\right]$
$\quad-4d\mathbb{E}\left[\left|\mathbb{E}_{i}\left[Y_{t_{i+1}}-\mathcal{Y}_{i+1}\right]\right|^{2}\right]-4d\mathbb{E}\left[\left|\mathbb{E}_{i}\left[\mathcal{Y}_{i+1}-\mathcal{NN}_{y,i+1,1+d,n}^{\varrho,L,\mathbf{n}}(p^{m}_{i+1},\theta^{*}_{y,i+1})\right]\right|^{2}\right]$
@@
$\mathbb{E}\left[\sum\limits_{i=0}^{N-1}\int_{t_{i}}^{t_{i+1}}|Z_{s}-\mathcal{V}^{m}_{t_{i}}|^{2}ds\right]$ $\leq\mathbb{E}\left[\sum\limits_{i=0}^{N-1}\int_{t_{i}}^{t_{i+1}}|Z_{s}-\bar{Z}_{t_{i}}|^{2}ds\right]+Ch+4d\mathbb{E}\left[\left|\Phi(X_{T})-\Phi(X_{N}^{\pi,m})\right|^{2}\right]$ (3.32)
@@
From (3.19) and (3.23), we deduce
$4d\mathbb{E}\left[\left|Y_{t_{i}}-\mathcal{Y}_{i}\right|^{2}\right]-4d\mathbb{E}\left[\left|\mathbb{E}_{i}\left[Y_{t_{i+1}}-\mathcal{Y}_{i+1}\right]\right|^{2}\right]$ (3.33)
@@
$\quad+\frac{16dL^{2}}{\gamma}\frac{1+\gamma h}{1-h}\left(Ch^{2}+\mathbb{E}\left[\int_{t_{i}}^{t_{i+1}}|Z_{s}-\mathcal{V}^{m}_{t_{i}}|^{2}ds\right]\right)+\frac{4d}{(1-h)h}\mathbb{E}\left[\left|\mathcal{Y}_{i}-\mathcal{U}^{m}_{t_{i}}\right|^{2}\right].$
Taking $\gamma=48dL^{2}$, we obtain $\frac{16dL^{2}}{\gamma}\frac{1+\gamma h}{1-h}\leq\frac{1}{2}$ for sufficiently small $h$. Then plugging (3.33) into (3.32), we have
$\frac{1}{2}\mathbb{E}\left[\sum\limits_{i=0}^{N-1}\int_{t_{i}}^{t_{i+1}}|Z_{s}-\mathcal{V}^{m}_{t_{i}}|^{2}ds\right]$ (3.34)
$\leq\mathbb{E}\left[\sum\limits_{i=0}^{N-1}\int_{t_{i}}^{t_{i+1}}|Z_{s}-\bar{Z}_{t_{i}}|^{2}ds\right]+Ch+C\max\limits_{0\leq i\leq N}\mathbb{E}\left[\left|Y_{t_{i}}-\mathcal{Y}_{i}\right|^{2}\right]+CN\sum\limits_{i=0}^{N-1}\mathbb{E}\left[\left|\mathcal{Y}_{i}-\mathcal{U}^{m}_{t_{i}}\right|^{2}\right]$

### 21 May 2026 v3

@@ 3.2 Error analysis
$|a+b|^{2}\leq(1+\gamma h)|a|^{2}+\left(1+\frac{1}{\gamma h}\right)|b|^{2},$
for $a,b$ in a Euclidean space and $\gamma>0$, together with the Lipschitz condition with respect to $f$, we deduce
$F_{z,i}^{r_{j}}(\theta_{z,i}^{r_{j}})=$ $\mathbb{E}\left[\left|\widetilde{\mathcal{V}}_{i}^{m}-\mathcal{NN}_{z,i,1+d,n\times d}^{\varrho,L,\mathbf{n}}(p_{i}^{m},\theta_{z,i}^{r_{j}})\right|^{2}\right].$ (35)
Moreover, by the optimality of $\theta_{z,i}^{*}$, we have, for any $\theta_{z,i}^{r_{j}}$,
$F_{z,i}^{*}(\theta_{z,i}^{*})$= $\mathbb{E}\left[\left|\widetilde{\mathcal{V}}_{i}^{m}-\mathcal{NN}_{z,i,1+d,n\times d}^{\varrho,L,\mathbf{n}}(p_{i}^{m},\theta_{z,i}^{*})\right|^{2}\right]$ (36)
$\leq$ $\ F_{z,i}^{r_{j}}(\theta_{z,i}^{r_{j}}).$
Furthermore,
$F_{y,i}^{r_{j}}(\theta_{y,i}^{r_{j}})=$ $\mathbb{E}\Bigg[\bigg|\widetilde{\mathcal{U}}_{i}^{m}+h\Big(f(t_{i},X_{i}^{\pi,m},\mathcal{NN}_{y,i,1+d,n}^{\varrho,L,\mathbf{n}}(p_{i}^{m},\theta_{y,i}^{r_{j}}),\mathcal{NN}_{z,i,1+d,n\times d}^{\varrho,L,\mathbf{n}}(p_{i}^{m},\theta_{z,i}^{*}))$ (37)
$\quad-f(t_{i},X_{i}^{\pi,m},\widetilde{\mathcal{U}}_{i}^{m},\widetilde{\mathcal{V}}_{i}^{m})\Big)-\mathcal{NN}_{y,i,1+d,n}^{\varrho,L,\mathbf{n}}(p_{i}^{m},\theta_{y,i}^{r_{j}})\bigg|^{2}\Bigg]$
$\leq$ $(1+Ch)\mathbb{E}\left[\left|\widetilde{\mathcal{U}}_{i}^{m}-\mathcal{NN}_{y,i,1+d,n}^{\varrho,L,\mathbf{n}}(p_{i}^{m},\theta_{y,i}^{r_{j}})\right|^{2}\right]$
$+Ch\mathbb{E}\left[\left|\widetilde{\mathcal{V}}_{i}^{m}-\mathcal{NN}_{z,i,1+d,n\times d}^{\varrho,L,\mathbf{n}}(p_{i}^{m},\theta_{z,i}^{*})\right|^{2}\right].$
Adding $hF_{z,i}^{*}(\theta_{z,i}^{*})$ to both sides of (37) and using (36), we have
$F_{y,i}^{r_{j}}(\theta_{y,i}^{r_{j}})+hF_{z,i}^{*}(\theta_{z,i}^{*})\leq$ $(1+Ch)\mathbb{E}\left[\left|\widetilde{\mathcal{U}}_{i}^{m}-\mathcal{NN}_{y,i,1+d,n}^{\varrho,L,\mathbf{n}}(p_{i}^{m},\theta_{y,i}^{r_{j}})\right|^{2}\right]$ (38)
$+Ch\mathbb{E}\left[\left|\widetilde{\mathcal{V}}_{i}^{m}-\mathcal{NN}_{z,i,1+d,n\times d}^{\varrho,L,\mathbf{n}}(p_{i}^{m},\theta_{z,i}^{r_{j}})\right|^{2}\right].$
On the other hand, by Young inequality in the form
$|a+b|^{2}\geq(1-\gamma h)|a|^{2}-\frac{1}{\gamma h}|b|^{2},$
for $a,b$ in a Euclidean space, $\gamma>0$, and $0<\gamma h<1$, we obtain
$F_{y,i}^{r_{j}}(\theta_{y,i}^{r_{j}})\geq$ $\left(1-\gamma h-\frac{2L^{2}h}{\gamma}\right)\mathbb{E}\left[\left|\widetilde{\mathcal{U}}_{i}^{m}-\mathcal{NN}_{y,i,1+d,n}^{\varrho,L,\mathbf{n}}(p_{i}^{m},\theta_{y,i}^{r_{j}})\right|^{2}\right]$ (39)
$-\frac{2L^{2}h}{\gamma}\mathbb{E}\left[\left|\widetilde{\mathcal{V}}_{i}^{m}-\mathcal{NN}_{z,i,1+d,n\times d}^{\varrho,L,\mathbf{n}}(p_{i}^{m},\theta_{z,i}^{*})\right|^{2}\right].$
Adding $hF_{z,i}^{*}(\theta_{z,i}^{*})$ to both sides of (39) gives
$F_{y,i}^{r_{j}}(\theta_{y,i}^{r_{j}})+hF_{z,i}^{*}(\theta_{z,i}^{*})$ (40)
$\geq$ $\left(1-\gamma h-\frac{2L^{2}h}{\gamma}\right)\mathbb{E}\left[\left|\widetilde{\mathcal{U}}_{i}^{m}-\mathcal{NN}_{y,i,1+d,n}^{\varrho,L,\mathbf{n}}(p_{i}^{m},\theta_{y,i}^{r_{j}})\right|^{2}\right]$
$+h\left(1-\frac{2L^{2}}{\gamma}\right)\mathbb{E}\left[\left|\widetilde{\mathcal{V}}_{i}^{m}-\mathcal{NN}_{z,i,1+d,n\times d}^{\varrho,L,\mathbf{n}}(p_{i}^{m},\theta_{z,i}^{*})\right|^{2}\right].$
Choosing $\gamma=4L^{2}$, we obtain, for $h$ small enough,
$F_{y,i}^{r_{j}}(\theta_{y,i}^{r_{j}})+hF_{z,i}^{*}(\theta_{z,i}^{*})\geq$ $(1-Ch)\mathbb{E}\left[\left|\widetilde{\mathcal{U}}_{i}^{m}-\mathcal{NN}_{y,i,1+d,n}^{\varrho,L,\mathbf{n}}(p_{i}^{m},\theta_{y,i}^{r_{j}})\right|^{2}\right]$ (41)
$+\frac{h}{2}\mathbb{E}\left[\left|\widetilde{\mathcal{V}}_{i}^{m}-\mathcal{NN}_{z,i,1+d,n\times d}^{\varrho,L,\mathbf{n}}(p_{i}^{m},\theta_{z,i}^{*})\right|^{2}\right].$
Thus, by the optimality of $\theta_{y,i}^{*}$, together with (38), we have
$(1-Ch)\mathbb{E}\left[\left|\widetilde{\mathcal{U}}_{i}^{m}-\mathcal{NN}_{y,i,1+d,n}^{\varrho,L,\mathbf{n}}(p_{i}^{m},\theta_{y,i}^{*})\right|^{2}\right]$ (42)
$+\frac{h}{2}\mathbb{E}\left[\left|\widetilde{\mathcal{V}}_{i}^{m}-\mathcal{NN}_{z,i,1+d,n\times d}^{\varrho,L,\mathbf{n}}(p_{i}^{m},\theta_{z,i}^{*})\right|^{2}\right]$
$\leq F_{y,i}^{*}(\theta_{y,i}^{*})+hF_{z,i}^{*}(\theta_{z,i}^{*})$
$\leq(1+Ch)\mathbb{E}\left[\left|\widetilde{\mathcal{U}}_{i}^{m}-\mathcal{NN}_{y,i,1+d,n}^{\varrho,L,\mathbf{n}}(p_{i}^{m},\theta_{y,i}^{r_{j}})\right|^{2}\right]$
$\quad+Ch\mathbb{E}\left[\left|\widetilde{\mathcal{V}}_{i}^{m}-\mathcal{NN}_{z,i,1+d,n\times d}^{\varrho,L,\mathbf{n}}(p_{i}^{m},\theta_{z,i}^{r_{j}})\right|^{2}\right].$
Taking the infimum over the neural network classes and using (25), we have, for $h$ small enough,
$\mathbb{E}\left[\left|\widetilde{\mathcal{U}}_{i}^{m}-\mathcal{NN}_{y,i,1+d,n}^{\varrho,L,\mathbf{n}}(p_{i}^{m},\theta_{y,i}^{*})\right|^{2}\right]+h\mathbb{E}\left[\left|\widetilde{\mathcal{V}}_{i}^{m}-\mathcal{NN}_{z,i,1+d,n\times d}^{\varrho,L,\mathbf{n}}(p_{i}^{m},\theta_{z,i}^{*})\right|^{2}\right]$ (43)
$\leq C\left(\widetilde{\varepsilon}_{i}^{y}+h\widetilde{\varepsilon}_{i}^{z}\right).$
@@
$\mathbb{E}\left[\left|\widetilde{\mathcal{V}}_{i}^{m}-\mathcal{V}_{i}^{m}\right|^{2}\right]\leq\frac{C}{K_{i}h}\leq\frac{C}{Kh},$ (44)
where $K=\min_{0\leq j\leq N-1}K_{j}$. The above inequality follows from the Monte Carlo approximation error together with the fact that the random variable
$\mathcal{NN}_{y,i+1,1+d,n}^{\varrho,L,\mathbf{n}}(p_{i+1}^{m,k},\theta_{y,i+1}^{*})\frac{(\Delta W_{i}^{m,k})^{\top}}{h}$
has a second moment of order $1/h$.
From (21), (23) and (44), Young inequality in the form
$|a+b|^{2}\leq(1+\gamma h)|a|^{2}+\left(1+\frac{1}{\gamma h}\right)|b|^{2},$
for $a,b$ in a Euclidean space and $\gamma>0$, together with the Lipschitz condition with respect to $f$, we obtain
@@
$\mathbb{E}\left[\left|\widetilde{\mathcal{U}}_{i}^{m}-\mathcal{U}_{i}^{m}\right|^{2}\right]$ $\leq(1+Ch)\mathbb{E}\Bigg[\bigg|\frac{1}{K_{i}}\sum\limits_{k=1}^{K_{i}}\mathcal{NN}_{y,i+1,1+d,n}^{\varrho,L,\mathbf{n}}(p_{i+1}^{m,k},\theta^{*}_{y,i+1})$ (45)
$\quad-\mathbb{E}_{i}\Big[\mathcal{NN}_{y,i+1,1+d,n}^{\varrho,L,\mathbf{n}}(p_{i+1}^{m},\theta^{*}_{y,i+1})\Big]\bigg|^{2}\Bigg]$
$\quad+Ch\mathbb{E}\left[\left|\widetilde{\mathcal{V}}_{i}^{m}-\mathcal{V}_{i}^{m}\right|^{2}\right]$
$\leq\frac{C}{K_{i}}\leq\frac{C}{K}.$
Plugging (43) and (45) into (34), we derive
$\mathbb{E}\left[\left|\mathcal{Y}_{i}-\mathcal{U}_{i}^{m}\right|^{2}\right]\leq$ $C\left(\widetilde{\varepsilon}_{i}^{y}+h\widetilde{\varepsilon}_{i}^{z}+\frac{1}{K}\right).$ (46)
From the definitions of $\widetilde{\varepsilon}_{i}^{y}$, $\varepsilon_{i}^{y}$ and Young inequality, we have
@@
$\widetilde{\varepsilon}_{i}^{y}$ $=\inf\limits_{\mathcal{NN}_{y,i,1+d,n}^{\varrho,L,\mathbf{n}}}\mathbb{E}\left[\left|\widetilde{\mathcal{U}}_{i}^{m}-\widetilde{u}(t_{i},X^{\pi,m}_{i})+\widetilde{u}(t_{i},X^{\pi,m}_{i})-\mathcal{NN}_{y,i,1+d,n}^{\varrho,L,\mathbf{n}}(p_{i}^{m},\theta^{r_{j}}_{y,i})\right|^{2}\right]$ (47)
$\leq 2\mathbb{E}\left[\left|\mathcal{U}_{i}^{m}-\widetilde{\mathcal{U}}_{i}^{m}\right|^{2}\right]+2\varepsilon_{i}^{y}.$
By (45), we restate (47) as
$\widetilde{\varepsilon}_{i}^{y}\leq C\left(\frac{1}{K_{i}}+\varepsilon_{i}^{y}\right)\leq C\left(\frac{1}{K}+\varepsilon_{i}^{y}\right).$ (48)
Similarly, from the definitions of $\widetilde{\varepsilon}_{i}^{z}$, $\varepsilon_{i}^{z}$, Young inequality and (44), we have
@@
$\widetilde{\varepsilon}_{i}^{z}$ $=\inf\limits_{\mathcal{NN}_{z,i,1+d,n\times d}^{\varrho,L,\mathbf{n}}}\mathbb{E}\left[\left|\widetilde{\mathcal{V}}_{i}^{m}-\widetilde{v}(t_{i},X^{\pi,m}_{i})+\widetilde{v}(t_{i},X^{\pi,m}_{i})-\mathcal{NN}_{z,i,1+d,n\times d}^{\varrho,L,\mathbf{n}}(p_{i}^{m},\theta_{z,i}^{r_{j}})\right|^{2}\right]$ (49)
$\leq 2\mathbb{E}\left[\left|\mathcal{V}_{i}^{m}-\widetilde{\mathcal{V}}_{i}^{m}\right|^{2}\right]+2\varepsilon_{i}^{z}$
$\leq C\left(\frac{1}{Kh}+\varepsilon_{i}^{z}\right).$
Plugging (48) and (49) into (46), we deduce
$\mathbb{E}\left[\left|\mathcal{Y}_{i}-\mathcal{U}_{i}^{m}\right|^{2}\right]\leq$ $C\left(\varepsilon_{i}^{y}+h\varepsilon_{i}^{z}+\frac{1}{K}\right).$ (50)
Inserting (50) into (33), we deduce
$\max\limits_{0\leq i\leq N-1}\mathbb{E}\left[\left|Y_{t_{i}}-\mathcal{Y}_{i}\right|^{2}\right]$ $\leq C\mathbb{E}\left[\left|g(X_{T})-g(X_{N}^{\pi,m})\right|^{2}\right]+Ch$ (51)
$\quad+C\mathbb{E}\left[\sum\limits_{i=0}^{N-1}\int_{t_{i}}^{t_{i+1}}|Z_{s}-\bar{Z}_{t_{i}}|^{2}ds\right]$
$\quad+C\sum\limits_{i=0}^{N-1}\left[N\varepsilon_{i}^{y}+\varepsilon_{i}^{z}+\frac{N}{K}\right].$
Step 3. Now, we prove the consistency of the $Z$-component. From the assumptions (ii)-(iii), Cauchy-Schwarz inequality and the tower property of conditional expectations, we have (see [20, 37])
$\mathbb{E}\left[\int_{t_{i}}^{t_{i+1}}|Z_{s}-\mathcal{V}^{m}_{i}|^{2}ds\right]$ (52)
@@
$=\mathbb{E}\left[\int_{t_{i}}^{t_{i+1}}|Z_{s}-\bar{Z}_{t_{i}}|^{2}ds\right]+2dh\mathbb{E}\left[\int_{t_{i}}^{t_{i+1}}f_{s}^{2}ds\right]$
$\quad+2d\mathbb{E}\left[\left|Y_{t_{i+1}}-\mathcal{Y}_{i+1}\right|^{2}\right]-2d\mathbb{E}\left[\left|\mathbb{E}_{i}\left[Y_{t_{i+1}}-\mathcal{Y}_{i+1}\right]\right|^{2}\right]$
@@
$\mathbb{E}\left[\sum\limits_{i=0}^{N-1}\int_{t_{i}}^{t_{i+1}}|Z_{s}-\mathcal{V}^{m}_{i}|^{2}ds\right]$ (53)
$\leq\mathbb{E}\left[\sum\limits_{i=0}^{N-1}\int_{t_{i}}^{t_{i+1}}|Z_{s}-\bar{Z}_{t_{i}}|^{2}ds\right]+Ch+4d\mathbb{E}\left[\left|g(X_{T})-g(X_{N}^{\pi,m})\right|^{2}\right]$
@@
From Young inequality in the form: $(a+b)^{2}\leq(1+\gamma h)a^{2}+(1+\frac{1}{\gamma h})b^{2},$ for $a,b\in\mathbb{R}$, and some $\gamma>0$ which will be given later, (31), Cauchy-Schwarz inequality, and the Lipschitz condition with respect to $f$, we deduce
$4d\mathbb{E}\left[\left|Y_{t_{i}}-\mathcal{Y}_{i}\right|^{2}\right]-4d\mathbb{E}\left[\left|\mathbb{E}_{i}\left[Y_{t_{i+1}}-\mathcal{Y}_{i+1}\right]\right|^{2}\right]$ (54)
@@
$\quad+\frac{16dL^{2}}{\gamma}\frac{1+\gamma h}{1-h}\left(Ch^{2}+\mathbb{E}\left[\int_{t_{i}}^{t_{i+1}}|Z_{s}-\mathcal{V}^{m}_{i}|^{2}ds\right]\right)+\frac{4d}{(1-h)h}\mathbb{E}\left[\left|\mathcal{Y}_{i}-\mathcal{U}^{m}_{i}\right|^{2}\right].$
Taking $\gamma=48dL^{2}$, we obtain $\frac{16dL^{2}}{\gamma}\frac{1+\gamma h}{1-h}\leq\frac{1}{2}$ for sufficiently small $h$. Then plugging (54) into (53), we have
$\frac{1}{2}\mathbb{E}\left[\sum\limits_{i=0}^{N-1}\int_{t_{i}}^{t_{i+1}}|Z_{s}-\mathcal{V}^{m}_{i}|^{2}ds\right]$ (55)
$\leq\mathbb{E}\left[\sum\limits_{i=0}^{N-1}\int_{t_{i}}^{t_{i+1}}|Z_{s}-\bar{Z}_{t_{i}}|^{2}ds\right]+Ch+C\max\limits_{0\leq i\leq N}\mathbb{E}\left[\left|Y_{t_{i}}-\mathcal{Y}_{i}\right|^{2}\right]+CN\sum\limits_{i=0}^{N-1}\mathbb{E}\left[\left|\mathcal{Y}_{i}-\mathcal{U}^{m}_{i}\right|^{2}\right]$